### v1

# Strategic inputs: feature selection from game-theoretic perspective

Abstract
The exponential growth of data volumes has led to escalating computational costs in machine learning model training. However, many features fail to contribute positively to model performance while consuming substantial computational resources. This paper presents an end-to-end feature selection framework for tabular data based on game theory. We formulate feature selection procedure based on a cooperative game where features are modeled as players, and their importance is determined through the evaluation of synergistic interactions and marginal contributions. The proposed framework comprises four core components: sample selection, game-theoretic feature importance evaluation, redundant feature elimination, and optimized model training. Experimental results demonstrate that the proposed method achieves substantial computation reduction while preserving predictive performance, thereby offering an efficient solution of the computational challenges of large-scale machine learning. The source code is available at https://github.com/vectorsss/strategy_inputs.
1 Introduction
A growth of data has created significant challenges for machine learning practitioners: it significantly increases computational cost and model performance. As datasets grow in both size and dimensionality, there is an actual need of creation of efficient feature selection methods that can reduce dimensionality while maintaining predictive performance. Feature selection is a crucial step in the machine learning pipeline since it directly impacts both model performance and training efficiency.
Development of explainable AI (XAI) has introduced new paradigms for feature selection based on feature importance evaluation. These XAI-based methods can be divided into two main groups: game-theoretic and non game-theoretic.
Non game-theoretic methods, such as Feature Importance (FI) (Breiman, 2001), Permutation Feature Importance (PFI) (Altmann et al., 2010), and Local Interpretable Model-Agnostic Explanations (LIME) (Ribeiro et al., 2016) are widely used in practice due to their computational efficiency and ease of implementation. However, these methods typically evaluate feature importance independently of other features, i.e., they cannot capture complex interactions between features.
Game-theoretic methods, particularly, those based on the Shapley value, provide a more comprehensive framework for feature importance evaluation. The SHAP method (SHapley Additive exPlanations) (Lundberg & Lee, 2017) provides a unified approach with various implementations including KernelSHAP, SamplingSHAP, etc. Recently developed ShapG method (Zhao et al., 2025a; b), inspired by Shapley value-based centrality measures on graphs (Zhao & Parilina, 2024), demonstrates improved computational efficiency compared to traditional Shapley-based methods. These game-theoretic approaches evaluate feature importance based on cooperative interactions among features, enabling more detailed understanding of feature contributions.
Despite their theoretical advantages, game-theoretic methods face significant computational challenges. The computational cost increases exponentially with the number of features when using the Shapley value to measure their importance. Both game-theoretic and non game-theoretic methods exhibit a linear relationship between data scale (number of observations) and computational cost for feature importance evaluation. Therefore, selecting a representative subset of observations can significantly reduce computational costs for feature importance evaluation. In the paper Axiotis et al. (2024), the authors proposed a clustering-based sensitivity sampling method to select diverse subsets of observations for efficient learning. In this work, we adopt a simpler clustering-based method that emphasizes balanced coverage of the data distribution, ensuring the sampled subset reflects both global and local diversity without over-emphasizing highly influential observations. The latter may lead to inaccurate feature importance estimators.
The main contributions of this paper are as follows:
  1. 1.
    We propose a novel clustering-based sampling method that ensures balanced (representative) coverage of the data distribution.
  2. 2.
    We move the feature importance evaluation from the training stage to the prediction stage, allowing efficient feature selection without retraining the model.
  3. 3.
    We propose a feature selection pipeline that integrates our sampling method and importance evaluation into a unified framework and benchmark its performance against state-of-the-art methods.
  4. 4.
    We demonstrate the effectiveness of the proposed method on various tabular datasets, showing that it accelerates the model training process without significant loss in model performance.
The rest of the paper is organized as follows: in Section 2 we describe the problem and the way how we are solving it. It includes data sampling procedure (Section 2.1), feature selection (Sections 2.2 and 2.3). Experiments on ten datasets are described in Section 3. We briefly conclude and present our future work in Section 4.
2 Problem formulation
In this section, we describe the problem of feature selection from a game-theoretic perspective and the proposed pipeline for this task.
We consider a supervised learning setting with dataset $D=\{({\bm{x}}^{(i)},y^{(i)})\}_{i=1}^{n}$ consisting of $n$ observations, where ${\bm{x}}^{(i)}=({x}_{1}^{(i)},\ldots,{x}_{M}^{(i)})\in\mathbb{R}^{M}$ represents the $i$th observation of a feature vector consisting of $M$ features, and $y^{(i)}\in\mathbb{R}$ is the $i$th observation of a target variable. We assume that the dataset is “cleaned” and preprocessed, meaning that there are no missing values, no outliers, etc. Otherwise, it is necessary to perform data preprocessing before feature selection.
There is a linear relationship between the data scale (number of observations) and the computational cost of feature importance evaluation. Therefore, we develop a two-stage feature selection pipeline, in which the first stage involves selecting a subset of representative observations from the original dataset (Section 2.1), and the second stage evaluates the feature importance based on the selected sample (Section 2.2).
2.1 Data sampling
We have dataset $D$ described above with $n$ observations and we aim in selecting a representative of sample size $s$, i.e., subset $S\subseteq D$ with $|S|=s<n$. The term “representative” means that the selected subset $S$ is able to express the original dataset $D$ in such a way that every observation from $D$ has a representative in $S$ within a bounded distance.
For this purpose, we propose a diversity-based sampling method, which includes the following steps:
  1. Step 1
    For a given $k$, find cluster centers $C=\{c_{1},\ldots,c_{k}\}$ (we use $k=\lfloor\sqrt{s}\rfloor$ as default under the assumption that data is balanced, i.e. in each cluster $C_{i}$ the number of observations is, roughly speaking, equal) minimizing the within-cluster variance:
    $C^{*}=\arg\min_{C}\sum_{x\in D}\min_{c\in C}\|x-c\|^{2}.$
    This can be solved by the $k$-means++ algorithm (see Lloyd (1982); Arthur & Vassilvitskii (2006)). By solving this minimization problem, we obtain the set of clusters $\{C_{1},\ldots,C_{k}\}$.
  2. Step 2
    For each cluster $C_{i}$, $i=1,\ldots,k$, select $s_{i}$ observations, where
    $s_{i}=\begin{cases}\lfloor\frac{s}{k}\rfloor+1,&\text{if }i\leq(s\mod k),\\ \lfloor\frac{s}{k}\rfloor,&\text{otherwise}.\end{cases}$
    Since we aim to select $s$ observations in total, we need to distribute them among $k$ clusters. Therefore, we need to select $\lfloor s/k\rfloor$ observations from each cluster $C_{i}$, $i=1,\ldots,k$, and then distribute the remaining $(s\mod k)$ observations among the first $(s\mod k)$ clusters. This ensures that the total number of observations selected is equal to $s$.
    For cluster $C_{i}$ with observations sorted by increasing distance to center $c_{i}$, let $\pi_{i}:\{1,\ldots,|C_{i}|\}\rightarrow C_{i}$ be a rank of the $i$-th point in cluster $C_{i}$, such that:
    $\|\pi_{i}(1)-c_{i}\|\leq\|\pi_{i}(2)-c_{i}\|\leq...\leq\|\pi_{i}(|C_{i}|)-c_{i}\|.$
    Select $s_{i}$ observations from $C_{i}$ with uniform intervals:
    $S_{i}=\{\pi_{i}(j\cdot\Delta_{i}):j\in\{0,1,...,s_{i}-1\}\text{ and }j\cdot\Delta_{i}<|C_{i}|\},$
    where $\Delta_{i}=\max(1,\lfloor|C_{i}|/s_{i}\rfloor)$ is the distance between selected observations.
This approach ensures both global diversity (through clustering) and local diversity (through evenly-spaced sampling within each cluster), providing a computationally efficient method that approximates maximizing the sum of pairwise distances in the selected subset.
2.2 Feature selection
Assume that we have a model $f:\mathbb{R}^{M}\rightarrow\mathbb{R}$, which receives an input vector ${\bm{x}}=({x}_{1},\ldots,{x}_{M})\in\mathbb{R}^{M}$, representing an observation of $M$ features, and gives the output $f({\bm{x}})$.
The feature selection problem can be formulated from a cooperative game perspective with the goal in identifying a subset of features that maximizes the predictive performance of a model. In other words, we can identify a subset of features that have low impact on the model performance and this subset should be removed from consideration. We formulate this problem using cooperative game theory, where each feature is treated as a player whose contributions to model performance can be measured.
A cooperative game is described by the pair $(\mathcal{M},v)$, where $\mathcal{M}=\{\mathcal{F}_{1},\ldots,\mathcal{F}_{M}\}$ represents the set of players/features, and $v:2^{\mathcal{M}}\rightarrow\mathbb{R}$ is a characteristic function assigning a “strength” to any coalition of players (i.e., any subset $\mathcal{S}\subset\mathcal{M}$, with $S=|\mathcal{S}|$). Specifically, the value $v(\mathcal{S})$ quantifies the “strength”, predictive power, or contribution of coalition $\mathcal{S}$.
The first step in feature selection procedure is to evaluate importance of all features, and the second step is to remove features with low importance. To evaluate $v(\mathcal{S})$ for any coalition, we define the hybrid input ${\bm{z}}^{(\mathcal{S})}=({z}_{1}^{(\mathcal{S})},\ldots,{z}_{M}^{(\mathcal{S})})\in\mathbb{R}^{M}$ by
${z}_{i}^{(\mathcal{S})}=\begin{cases}{x}_{i},&\text{if }i\in\mathcal{S},\\ r_{i},&\text{if }i\notin\mathcal{S}.\end{cases}$ (1)
If a feature belongs to coalition $\mathcal{S}$, its values from dataset $D$ are collected. Otherwise, the reference value of a feature is calculated based on dataset $D$. Given dataset $D$, we can obtain a reference value $r_{i}$ for each feature $\mathcal{F}_{i}$, calculated as a statistic from the feature observations (such as mean, median, or mode). This hybrid input ${z}_{i}^{(\mathcal{S})}$ replaces features absent in coalition $\mathcal{S}$ with their reference values while preserving the values of features in $\mathcal{S}$.
There are several methods to define characteristic function $v(\mathcal{S})$:
  1. 1.
    Method 1. Model output:
    $v(\mathcal{S})=f({\bm{z}}^{(\mathcal{S})}),$ (2)
    which is the most straightforward approach used in many XAI methods based on cooperative game theory. The change in model output when a feature is present versus it is absent indicates a feature importance. If the output does not change, the feature is unimportant. The magnitude of change reflects the degree of importance.
  2. 2.
    Method 2. Sample-level performance-based metric: characteristic function $v(\mathcal{S})$ is defined as follows:
[…]
    where $\mathbf{1}_{\{\cdot\}}$ is the indicator function. For classification, this returns +1+1 for correct predictions and 1-1 for incorrect ones. Negative values indicate features that harm model performance.
  3. 3.
    Method 3. Global-model performance: Similar to equation 3, but using dataset-wide metrics: $R^{2}$, negative MAE, or negative RMSE for regression tasks; accuracy or F1 score for classification tasks (Zhao et al. (2025a)).
The first two methods provide sample-level metrics, while the third provides global metrics. For sample-level metrics, we can obtain a global feature importance by averaging across all samples. When using model output (Method 1), we average the absolute values since both increases and decreases indicate importance. Global performance metrics (Method 3) directly yields feature importance scores without averaging.
Once the values of characteristic function $v(\mathcal{S})$ are defined using one of the proposed methods, a challenge in cooperative game theory is in determining a “fair” allocation of the total payoff $v(\mathcal{M})$ (the value of the grand coalition) distributing its payoff among participating players.
The Shapley value (Shapley, 1951) is a widely used solution of this problem:
$\phi_{i}=\sum_{\mathcal{S}\subseteq\mathcal{M}\setminus\{i\}}\frac{|\mathcal{S}|!(M-|\mathcal{S}|-1)!}{M!}\left(v(\mathcal{S}\cup\{i\})-v(\mathcal{S})\right),i\in\mathcal{M}.$
The Centre of the Imputation Set (CIS) Value (Driessen & Funaki, 1991) offers a more computationally efficient alternative:
$CIS_{i}(v)=v(\{i\})+M^{-1}\left[v(\mathcal{M})-\sum_{j\in\mathcal{M}}v(\{j\})\right],i\in\mathcal{M},$
and it requires computation of only $M+1$ values of $v$ compared to $2^{M}$ values for the exact Shapley value.
To further improve computational efficiency, we adapt the graph-based sampling approach from the ShapG method (Zhao et al., 2025a). ShapG constructs a collaborator set $\mathcal{C}_{i}$ for each feature $i$, restricting feature interactions to within these sets during evaluation. When this set is not large, i.e. $|\mathcal{C}_{i}|<\ell$ (default $\ell=15$), the exact Shapley value can be computed; otherwise, Monte Carlo approximation is used. We integrate this approach into our pipeline without requiring model retraining, significantly reducing computational costs.
In this paper, we focus primarily on the sample-level performance-based metric for feature importance evaluation since it provides fine-grained insights into feature contributions. Global model performance metrics are used as a post-processing step to validate the overall effectiveness of the selected features.
2.3 Proposed Pipeline for Feature Selection
Our feature selection pipeline consists of the following steps as shown in Figure 1.
  1. 1.
    Data Preprocessing [Optional]: If dataset contains outliers, apply appropriate outlier detection and removal techniques. Otherwise, proceed with the next step.
  2. 2.
    Initial Model Training: Train an initial model on a representative subset for a limited number of epochs (e.g., 5-10 epochs) to obtain a preliminary model.
  3. 3.
    Diversity Sampling: For large datasets (e.g., $n>10^{4}$ observations), apply the diversity sampling method described in Section 2.1 to select a representative subset of observations for feature importance evaluation.
  4. 4.
    Feature Importance Evaluation: Using the trained initial model, compute feature importance scores for all features using the characteristic functions defined in Section 2.2 (e.g., model output or negative squared error).
  5. 5.
    Feature Selection: Remove features with importance scores below a threshold $\tau$, or only keep the top $q\%$ of features. Specifically, retain feature $i$ if $\phi_{i}>\tau$ or $\phi_{i}$ is in the top $q\%$ of overall importance scores, where $\phi_{i}$ is the Shapley value or CIS value $i$th component.
  6. 6.
    Model training: Train a final model using only the selected features on the complete dataset (or the sampled subset for very large datasets).
DatasetContainsOutliers?DataPreprocessingInitial ModelTraining(5-10 epochs)LargeDataset?($n>10^{4}$)DiversitySamplingUse FullDatasetFeature ImportanceEvaluation(Shapley/CIS)Feature Selection($\phi_{i}>\tau$ or top $q\%$)Model Trainingwith SelectedFeaturesFinal ModelYesNoYesNo123456OptionalRequiredOutputLegend:
Figure 1: Proposed pipeline for feature selection.
3 Experiments
In this section, we evaluate how our proposed feature selection pipeline affects the performance of deep-learning models on tabular data, focusing on MLP and TabM (Gorishniy et al., 2025), the current state-of-the-art on TabRed benchmarks (Rubachev et al., 2024).
3.1 Experimental setup
3.1.1 Models
We evaluate four model configurations based on the TabM benchmark suite:
  1. 1.
    MLP: A multilayer perceptron with fully connected layers [^1] We use the same architecture as in Gorishniy et al. (2025), the number of layers varies according to the dataset., ReLU activations, and task-specific output heads.
  2. 2.
    TabM: A parameter-efficient ensemble model for tabular data that trains multiple MLPs simultaneously while sharing most parameters between them, achieving state-of-the-art performance (Gorishniy et al., 2025).
  3. 3.
    MLP+PLE: MLP enhanced with PLE (Piecewise Linear Embedding, see Gorishniy et al. (2022)) for numerical features, which splits continuous values into segments and learns separate representations for each segment, enabling better modeling of non-linear relationships.
  4. 4.
    TabM+PLE: TabM with PLE, combining attention mechanisms with sophisticated numerical encoding.
3.1.2 Configuration
Table 1: Experiment Configuration
Parameter Value
Feature importance method ShapG
Characteristic function Sample-level performance (Eq. 3)
Feature retention (top-$q\%$) 80%
Sampling threshold $\ell$ 5
Number of seeds 15
Training strategy Early stopping (patience=16)
Distance metric for collaboration set $\mathcal{C}_{i}$ construction 1 - cosine similarity
Evaluation metrics Accuracy (classification), $R^{2}$ (regression)
We compare baseline models (all the features included) against models with feature selection (FS), reporting mean $\pm$ std of model performance and training time across 15 seeds.
3.2 Datasets
In total, we use 10 datasets from (Gorishniy et al., 2025).
Table 2: Dataset configuration from Gorishniy et al. (2025)
Name Abbr #Train #Validation #Test #Num #Bin #Cat Task type Batch size
Adult AD 26 048 6 513 16 281 6 1 8 Binclass 256
Black Friday BF 106 764 26 692 33 365 4 1 4 Regression 512
California Housing CH 13 209 3 303 4 128 8 0 0 Regression 256
Churn Modelling CM 6 400 1 600 2 000 7 3 1 Binclass 128
Covertype CO 371 847 92 962 116 203 10 4 1 Multiclass 1024
Diamond DI 34 521 8 631 10 788 6 0 3 Regression 512
Higgs Small HS 62 751 15 688 19 610 28 0 0 Binclass 512
House 16H HO 14 581 3 646 4 557 16 0 0 Regression 256
Microsoft MI 723 412 235 259 241 521 131 5 0 Regression 1024
Otto Group Products OT 39 601 9 901 12 376 93 0 0 Multiclass 512
3.3 Implementation details
Training. We use the TabM benchmark suite with hyperparameters from Gorishniy et al. (2025), tuned via Optuna (Akiba et al., 2019).
Feature importance. We adapt ShapG (Zhao et al., 2025a; b) by moving evaluation from training to prediction stage, avoiding model retraining for each feature subset. Following Zhao et al. (2025b)’s improved strategy, we construct collaboration sets $\mathcal{C}_{i}$ through a graph-based approach: (1) rank features by their correlation with the target using cosine distance (1 - cosine similarity), (2) build a complete feature graph with edge weights based on pairwise cosine distance, (3) prune edges starting from the least correlated features (processed in ascending order of importance) while maintaining graph connectivity, targeting a graph density of $1.5\times$ the minimum connectivity ratio (capped at 1.0). This approach better captures feature interactions compared to the original method in Zhao et al. (2025a).
3.4 Results & analysis
Under the experimental setup described above, we evaluate performance of the proposed feature selection pipeline on 10 benchmark datasets. Our analysis reveals distinct patterns between classification and regression tasks, as shown in Figure 2. On the $OX$-axis of each graph, the four models (MLP, TabM, MLP+PLE, TabM+PLE) are presented. Ten subfigures correspond to ten datasets presented in Table 2. Accuracy (for classification tasks) or $R^{2}$ (for regression tasks) of the model is presented on the $OY$-axis. The purple (green) column corresponds to the baseline (top-80% feature) method.
Figure 2: Comparison of model performances across 10 datasets.
Classification tasks. Classification tasks demonstrate remarkable robustness to feature reduction, with accuracy degradation ranging from 0.5% to 8.7% and average 3% across all model-dataset combinations. We notice that the adult, otto, and covtype2 datasets maintain performance particularly well, with average accuracy drops below 2%. The most significant degradation occurs in the churn and higgs-small datasets. Our explanation of such a drop is that these tasks rely more significantly on the complete set of features.
Regression tasks. In contrast to classification, regression tasks exhibit substantially larger performance degradation under feature selection, with $R^{2}$ reductions ranging from 3.9% to 17.9%. The california and house datasets show the most pronounced sensitivity to feature reduction, with average $R^{2}$ drops exceeding 10%. The diamond dataset demonstrates relative stability, maintaining over 95% of its baseline $R^{2}$ performance. The potential reason for this sensitivity to reduction of features could be attributed to the bad quality of baseline models ($R^{2}$ is close to 1 for the diamond dataset, but it is significantly lower for other datasets).
Training time. Figure 3 presents the training time (presented on the $OY$-axis) comparison between baseline (purple columns) and top-80% feature (green columns) configurations. We observe that the training time for the top-80% feature configurations is not always lower than for the baseline ones. This can be explained by the fact that in our experiments, we did not set the limitation for epochs, we only have set the patience for early stopping. Therefore, the training time depends on the convergence of the model, in some cases, the model with fewer features converges faster, while in other cases, it takes longer time to converge.
Figure 3: Comparison of training time of different methods across 10 datasets.
Table 3: Summary of results: Baseline vs Top 80% Feature Selection
Model Classification (Accuracy) Regression ($R^{2}$) Overall Time
Baseline Top 80% $\Delta$% Baseline Top 80% $\Delta$% $\Delta$% $\Delta$%
MLP 0.8423 0.8157 -3.2 0.6327 0.5824 -7.9 -5.2 -7.2
MLP-PLE 0.8458 0.8257 -2.4 0.6448 0.5762 -10.6 -6.0 +1.3
TabM 0.8520 0.8254 -3.1 0.6507 0.5739 -11.8 -6.9 -19.9
TabM-PLE 0.8557 0.8264 -3.4 0.6568 0.6050 -7.9 -5.4 -23.2
Summary. Table 3 summarizes the results of our experiments. Feature selection successfully reduces the number of features while maintaining model performance comparable to using all features, with an average drop of 3% in classification accuracy, 9.6% in $R^{2}$ in regression task, and 5.9% overall. Additionally, training time is reduced by in average of 12.3%, though the effect varies with the model: MLP-PLE becomes slightly slower (+1.3%), MLP shows a modest speed-up (-7.2%), and both TabM and TabM-PLE achieve substantial reduction of time (-19.9% and -23.2%, respectively). These results allow to conclude that feature selection offers a favorable trade-off between predictive performance and computational efficiency, with particular benefits for more complex model architectures.
4 Conclusion and future work
In this work, we proposed a feature selection pipeline based on game-theoretic feature importance evaluation. Our pipeline employs a k-means based diversity sampling procedure to select a representative subset of observations for feature importance evaluation, thereby reducing the influence of outliers and ensuring more robust feature ranking. We demonstrated the effectiveness of this approach on 10 benchmark datasets using state-of-the-art tabular deep learning models. The results show that integrating our feature selection pipeline into model training reduces computational costs by 12.3% on average while maintaining competitive performance, with accuracy degradation by 5.9% on average across all tasks.
Our study has several limitations that open opportunities for future research. First, we evaluated only a single feature selection configuration (retaining the top 80% of features by importance) using the ShapG method, which showed effectiveness in recent feature importance evaluation studies Zhao et al. (2025a; b). Comprehensive comparisons with alternative feature selection approaches—including LIME, Permutation Feature Importance, and other game-theoretic methods with different characteristic functions—remain to be explored. We can also extend the proposed method by varying the number of features selected (the percentage of $q$-top features) for a fixed drop in the model performance. This extension is due to varying the number of significant features depending of the dataset and deep-learning model using as a basis.
Additionally, extending this benchmark to industrial datasets, such as those described in Rubachev et al. (2024), would provide valuable insights on performance of real-world data analysis.
The proposed diversity sampling method also shows promising results beyond feature importance evaluation. Preliminary experiments suggest its effectiveness for training data sampling, in particular, for domains with heterogeneous data distributions. While we have observed encouraging results on proprietary wireless communication datasets with TabM (Gorishniy et al., 2025) and FT-transformer (Gorishniy et al., 2021) models, we cannot present these findings due to confidentiality constraints. Future work should investigate this application systematically on publicly available datasets.
These findings suggest that feature selection remains a valuable preprocessing step even for modern deep learning approaches to tabular data, offering a practical way to data analysts to balance computational efficiency with respect to predictive performance.

### v2

@@
 These XAI-based methods can be divided into two main groups: game-theoretic and non game-theoretic.
-Non game-theoretic methods, such as Feature Importance (FI) (Breiman, 2001), Permutation Feature Importance (PFI) (Altmann et al., 2010), and Local Interpretable Model-Agnostic Explanations (LIME) (Ribeiro et al., 2016) are widely used in practice due to their computational efficiency and ease of implementation.
+Non game-theoretic methods, such as Feature Importance (FI) (breiman2001random), Permutation Feature Importance (PFI) (altmann2010permutation), and Local Interpretable Model-Agnostic Explanations (LIME) (ribeiro2016should) are widely used in practice due to their computational efficiency and ease of implementation.
 However, these methods typically evaluate feature importance independently of other features, i.e., they cannot capture complex interactions between features.
 Game-theoretic methods, particularly, those based on the Shapley value, provide a more comprehensive framework for feature importance evaluation.
-The SHAP method (SHapley Additive exPlanations) (Lundberg & Lee, 2017) provides a unified approach with various implementations including KernelSHAP, SamplingSHAP, etc.
-Recently developed ShapG method (Zhao et al., 2025a; b), inspired by Shapley value-based centrality measures on graphs (Zhao & Parilina, 2024), demonstrates improved computational efficiency compared to traditional Shapley-based methods.
+The SHAP method (SHapley Additive exPlanations) (lundberg2017unified) provides a unified approach with various implementations including KernelSHAP, SamplingSHAP, etc.
+Recently developed ShapG method (zhao2024shapgnewfeatureimportance; zhao2025complete), inspired by Shapley value-based centrality measures on graphs (zhao2024centralitymeasuresopiniondynamics), demonstrates improved computational efficiency compared to traditional Shapley-based methods.
+A sparsity-based deterministic method for the Shapley value approximation presented in erofeeva2025novel introduces a novel compressed-sensing perspective on this problem.
+However, as the number of players (features) increases, the associated time and memory requirements grow rapidly, which may limit its applicability in high-dimensional settings.
 These game-theoretic approaches evaluate feature importance based on cooperative interactions among features, enabling more detailed understanding of feature contributions.
+A detailed description and benchmark results for each method can be found in (zhao2025shapley).
 Despite their theoretical advantages, game-theoretic methods face significant computational challenges.
@@
 Therefore, selecting a representative subset of observations can significantly reduce computational costs for feature importance evaluation.
-In the paper Axiotis et al.
-(2024), the authors proposed a clustering-based sensitivity sampling method to select diverse subsets of observations for efficient learning.
+In the paper axiotis2024data, the authors proposed a clustering-based sensitivity sampling method to select diverse subsets of observations for efficient learning.
 In this work, we adopt a simpler clustering-based method that emphasizes balanced coverage of the data distribution, ensuring the sampled subset reflects both global and local diversity without over-emphasizing highly influential observations.
@@
     $C^{*}=\arg\min_{C}\sum_{x\in D}\min_{c\in C}\|x-c\|^{2}.$
-    This can be solved by the $k$-means++ algorithm (see Lloyd (1982); Arthur & Vassilvitskii (2006)). By solving this minimization problem, we obtain the set of clusters $\{C_{1},\ldots,C_{k}\}$.
+    This can be solved by the $k$-means++ algorithm (see lloyd1982least; arthur2006k). By solving this minimization problem, we obtain the set of clusters $\{C_{1},\ldots,C_{k}\}$.
   2. Step 2
@@
     Method 3.
-Global-model performance: Similar to equation 3, but using dataset-wide metrics: $R^{2}$, negative MAE, or negative RMSE for regression tasks; accuracy or F1 score for classification tasks (Zhao et al.
-(2025a)).
+Global-model performance: Similar to equation 3, but using dataset-wide metrics: $R^{2}$, negative MAE, or negative RMSE for regression tasks; accuracy or F1 score for classification tasks (zhao2024shapgnewfeatureimportance).
 The first two methods provide sample-level metrics, while the third provides global metrics.
@@
 Once the values of characteristic function $v(\mathcal{S})$ are defined using one of the proposed methods, a challenge in cooperative game theory is in determining a “fair” allocation of the total payoff $v(\mathcal{M})$ (the value of the grand coalition) distributing its payoff among participating players.
-The Shapley value (Shapley, 1951) is a widely used solution of this problem:
+The Shapley value (shapley1951notes) is a widely used solution of this problem:
 $\phi_{i}=\sum_{\mathcal{S}\subseteq\mathcal{M}\setminus\{i\}}\frac{|\mathcal{S}|!(M-|\mathcal{S}|-1)!}{M!}\left(v(\mathcal{S}\cup\{i\})-v(\mathcal{S})\right),i\in\mathcal{M}.$
-The Centre of the Imputation Set (CIS) Value (Driessen & Funaki, 1991) offers a more computationally efficient alternative:
+The Centre of the Imputation Set (CIS) Value (driessen1991coincidence) offers a more computationally efficient alternative:
 $CIS_{i}(v)=v(\{i\})+M^{-1}\left[v(\mathcal{M})-\sum_{j\in\mathcal{M}}v(\{j\})\right],i\in\mathcal{M},$
 and it requires computation of only $M+1$ values of $v$ compared to $2^{M}$ values for the exact Shapley value.
-To further improve computational efficiency, we adapt the graph-based sampling approach from the ShapG method (Zhao et al., 2025a).
+To further improve computational efficiency, we adapt the graph-based sampling approach from the ShapG method (zhao2024shapgnewfeatureimportance).
 ShapG constructs a collaborator set $\mathcal{C}_{i}$ for each feature $i$, restricting feature interactions to within these sets during evaluation.
@@
 3 Experiments
-In this section, we evaluate how our proposed feature selection pipeline affects the performance of deep-learning models on tabular data, focusing on MLP and TabM (Gorishniy et al., 2025), the current state-of-the-art on TabRed benchmarks (Rubachev et al., 2024).
+In this section, we evaluate how our proposed feature selection pipeline affects the performance of deep-learning models on tabular data, focusing on MLP and TabM (gorishniy2024tabm), the current state-of-the-art on TabRed benchmarks (rubachev2024tabred).
 3.1 Experimental setup
@@
   1. 1.
-    MLP: A multilayer perceptron with fully connected layers [^1] We use the same architecture as in Gorishniy et al.
-(2025), the number of layers varies according to the dataset., ReLU activations, and task-specific output heads.
+    MLP: A multilayer perceptron with fully connected layers [^1] We use the same architecture as in gorishniy2024tabm, the number of layers varies according to the dataset., ReLU activations, and task-specific output heads.
   2. 2.
-    TabM: A parameter-efficient ensemble model for tabular data that trains multiple MLPs simultaneously while sharing most parameters between them, achieving state-of-the-art performance (Gorishniy et al., 2025).
+    TabM: A parameter-efficient ensemble model for tabular data that trains multiple MLPs simultaneously while sharing most parameters between them, achieving state-of-the-art performance (gorishniy2024tabm).
   3. 3.
-    MLP+PLE: MLP enhanced with PLE (Piecewise Linear Embedding, see Gorishniy et al.
-(2022)) for numerical features, which splits continuous values into segments and learns separate representations for each segment, enabling better modeling of non-linear relationships.
+    MLP+PLE: MLP enhanced with PLE (Piecewise Linear Embedding, see gorishniy2022embeddings) for numerical features, which splits continuous values into segments and learns separate representations for each segment, enabling better modeling of non-linear relationships.
   4. 4.
@@
 Parameter Value
-Feature importance method ShapG
+Feature importance method ShapG and PFI
 Characteristic function Sample-level performance (Eq. 3)
@@
 Sampling threshold $\ell$ 5
-Number of seeds 15
+Number of seeds 3
 Training strategy Early stopping (patience=16)
@@
 Evaluation metrics Accuracy (classification), $R^{2}$ (regression)
-We compare baseline models (all the features included) against models with feature selection (FS), reporting mean $\pm$ std of model performance and training time across 15 seeds.
+We compare baseline models (all the features included) against models with feature selection (FS), reporting mean $\pm$ std of model performance and training time across 3 seeds.
 3.2 Datasets
-In total, we use 10 datasets from (Gorishniy et al., 2025).
-Table 2: Dataset configuration from Gorishniy et al. (2025)
+In total, we use 10 datasets from (gorishniy2024tabm).
+Table 2: Dataset configuration from gorishniy2024tabm
 Name Abbr #Train #Validation #Test #Num #Bin #Cat Task type Batch size
@@
 3.3 Implementation details
-Training. We use the TabM benchmark suite with hyperparameters from Gorishniy et al. (2025), tuned via Optuna (Akiba et al., 2019).
+Training. We use the TabM benchmark suite with hyperparameters from gorishniy2024tabm, tuned via Optuna (akiba2019optuna).
 Feature importance.
-We adapt ShapG (Zhao et al., 2025a; b) by moving evaluation from training to prediction stage, avoiding model retraining for each feature subset.
-Following Zhao et al.
-(2025b)’s improved strategy, we construct collaboration sets $\mathcal{C}_{i}$ through a graph-based approach: (1) rank features by their correlation with the target using cosine distance (1 - cosine similarity), (2) build a complete feature graph with edge weights based on pairwise cosine distance, (3) prune edges starting from the least correlated features (processed in ascending order of importance) while maintaining graph connectivity, targeting a graph density of $1.5\times$ the minimum connectivity ratio (capped at 1.0).
-This approach better captures feature interactions compared to the original method in Zhao et al.
-(2025a).
+Since many datasets in our benchmark have more than 20 features, most existing feature-importance XAI methods are nearly infeasible in this setting.
+Thus, we focus on two methods for feature importance evaluation:
+  1. 1.
+    ShapG method: We adapt ShapG (zhao2024shapgnewfeatureimportance; zhao2025complete) by moving evaluation from training to prediction stage, avoiding model retraining for each feature subset.
+Following zhao2025complete’s improved strategy, we construct collaboration sets $\mathcal{C}_{i}$ through a graph-based approach: (1) rank features by their correlation with the target using cosine distance (1 - cosine similarity), (2) build a complete feature graph with edge weights based on pairwise cosine distance, (3) prune edges starting from the least correlated features (processed in ascending order of importance) while maintaining graph connectivity, targeting a graph density of $1.5\times$ the minimum connectivity ratio (capped at 1.0).
+This approach better captures feature interactions compared to the original method in zhao2024shapgnewfeatureimportance.
+  2. 2.
+    PFI method: We implement the Permutation Feature Importance (PFI) method (altmann2010permutation) by measuring the drop in a model performance when each feature values are randomly shuffled across samples.
+The algorithm shuffles each feature several times (5 by default), measures the effect after each shuffle, and averages the results to reduce randomness.
+Features giving larger drop in the performance when shuffled are deemed more important, providing a model-agnostic evaluation of feature importance without requiring model retraining.
 3.4 Results & analysis
@@
 Accuracy (for classification tasks) or $R^{2}$ (for regression tasks) of the model is presented on the $OY$-axis.
-The purple (green) column corresponds to the baseline (top-80% feature) method.
-Figure 2: Comparison of model performances across 10 datasets.
+The purple columns correspond to the baseline without feature selection, while the green and light-blue columns correspond to models using ShapG and Permutation feature selection (top 80% features), respectively.
+Figure 2: Model performance ($R^{2}$ for regression task and accuracy for classification task) across 10 datasets.
 Classification tasks.
-Classification tasks demonstrate remarkable robustness to feature reduction, with accuracy degradation ranging from 0.5% to 8.7% and average 3% across all model-dataset combinations.
-We notice that the adult, otto, and covtype2 datasets maintain performance particularly well, with average accuracy drops below 2%.
-The most significant degradation occurs in the churn and higgs-small datasets.
-Our explanation of such a drop is that these tasks rely more significantly on the complete set of features.
+For classification tasks (datasets Adult, Churn Modelling, Covertype, Higgs Small, Otto Group Products), both feature-selection methods (ShapG and PFI) show consistently strong robustness to feature reduction.
+Across all datasets, both ShapG and Permutation selection of the top 80% most important features yield accuracy that remains close to the baseline models trained on all features (purple bars), with typical degradation of only a few percentage points and an average drop less than 4% across all model-dataset combinations.
+Notably, on datasets such as Adult, both ShapG and PFI maintain performance very close to the baseline.
+In some cases (datasets Covertype and Otto Group Products), ShapG significantly outperforms PFI in terms of average accuracy, whereas on the Higgs Small dataset, PFI substantially outperforms ShapG.
+For the Churn Modelling dataset, both methods exhibit comparable performance.
 Regression tasks.
-In contrast to classification, regression tasks exhibit substantially larger performance degradation under feature selection, with $R^{2}$ reductions ranging from 3.9% to 17.9%.
-The california and house datasets show the most pronounced sensitivity to feature reduction, with average $R^{2}$ drops exceeding 10%.
-The diamond dataset demonstrates relative stability, maintaining over 95% of its baseline $R^{2}$ performance.
-The potential reason for this sensitivity to reduction of features could be attributed to the bad quality of baseline models ($R^{2}$ is close to 1 for the diamond dataset, but it is significantly lower for other datasets).
+In regression tasks (datasets Black Friday, California Housing, House 16H, Diamond, Microsoft), feature reduction affects the two methods very differently: ShapG shows substantial sensitivity, while the PFI method maintains stable performance.
+Across all regression datasets, ShapG exhibits substantial degradation in $R^{2}$, often exceeding 10%, that is observed in datasets California Housing and House 16H.
+In contrast, the PFI method remains much more stable: its performance drops are consistently small, typically below 6%, and in some cases even slightly surpass or remain close to the baseline (e.g., datasets California Housing and Diamond).
+This suggests that for regression tasks, the PFI method provides a more reliable feature importance evaluation compared to ShapG.
 Training time.
-Figure 3 presents the training time (presented on the $OY$-axis) comparison between baseline (purple columns) and top-80% feature (green columns) configurations.
-We observe that the training time for the top-80% feature configurations is not always lower than for the baseline ones.
-This can be explained by the fact that in our experiments, we did not set the limitation for epochs, we only have set the patience for early stopping.
-Therefore, the training time depends on the convergence of the model, in some cases, the model with fewer features converges faster, while in other cases, it takes longer time to converge.
-Figure 3: Comparison of training time of different methods across 10 datasets.
-Table 3: Summary of results: Baseline vs Top 80% Feature Selection
-Model Classification (Accuracy) Regression ($R^{2}$) Overall Time
-Baseline Top 80% $\Delta$% Baseline Top 80% $\Delta$% $\Delta$% $\Delta$%
-MLP 0.8423 0.8157 -3.2 0.6327 0.5824 -7.9 -5.2 -7.2
-MLP-PLE 0.8458 0.8257 -2.4 0.6448 0.5762 -10.6 -6.0 +1.3
-TabM 0.8520 0.8254 -3.1 0.6507 0.5739 -11.8 -6.9 -19.9
-TabM-PLE 0.8557 0.8264 -3.4 0.6568 0.6050 -7.9 -5.4 -23.2
+As shown in Figure 3, the impact of feature selection on training time varies across models and datasets.
+Notably, the 80% of feature configurations of both ShapG and PFI do not always lead to smaller training time compared with the baseline.
+This arises because the number of training epochs is not fixed in our setting; instead, we rely on early stopping with a patience parameter.
+Consequently, the actual training time is determined by the convergence dynamics of each model.
+In some cases, reducing the feature set accelerates convergence, whereas in others it may slow it down, leading to longer training time.
+Figure 3: Training time for different methods across 10 datasets.
+Table 3: Performance across different feature selection methods
+XAI Method Model Classification Regression
+Accuracy Time (s) $R^{2}$ Time (s)
+Baseline MLP 0.8426 22.1 0.6329 16.9
+MLP-PLE 0.8466 27.5 0.6443 24.7
+TabM 0.8520 326.6 0.6510 277.1
+TabM-PLE 0.8554 345.7 0.6561 328.8
+ShapG MLP 0.8182 (-3.1%) 22.2 (-6.6%) 0.5865 (-7.1%) 13.9 (-8.4%)
+MLP-PLE 0.8226 (-3.0%) 26.3 (+4.3%) 0.5750 (-10.1%) 25.0 (-8.1%)
+TabM 0.8153 (-4.3%) 199.9 (-10.8%) 0.5760 (-11.3%) 318.9 (-20.3%)
+TabM-PLE 0.8284 (-3.4%) 353.0 (-21.5%) 0.6085 (-7.4%) 212.2 (-38.8%)
+Avg. -3.5% -8.7% -9.0% -18.9%
+PFI MLP 0.8077 (-4.1%) 25.8 (+4.2%) 0.6262 (-1.2%) 17.2 (+4.8%)
+MLP-PLE 0.8214 (-3.0%) 31.7 (+7.2%) 0.6393 (-0.9%) 27.2 (+6.9%)
+TabM 0.8237 (-3.4%) 338.4 (-10.0%) 0.6392 (-1.8%) 415.0 (-6.7%)
+TabM-PLE 0.8197 (-4.3%) 548.2 (-21.8%) 0.6451 (-1.8%) 299.6 (-21.7%)
+Avg. -3.7% -5.1% -1.4% -4.2%
 Summary.
-Table 3 summarizes the results of our experiments.
-Feature selection successfully reduces the number of features while maintaining model performance comparable to using all features, with an average drop of 3% in classification accuracy, 9.6% in $R^{2}$ in regression task, and 5.9% overall.
-Additionally, training time is reduced by in average of 12.3%, though the effect varies with the model: MLP-PLE becomes slightly slower (+1.3%), MLP shows a modest speed-up (-7.2%), and both TabM and TabM-PLE achieve substantial reduction of time (-19.9% and -23.2%, respectively).
-These results allow to conclude that feature selection offers a favorable trade-off between predictive performance and computational efficiency, with particular benefits for more complex model architectures.
-4 Conclusion and future work
+Table 3 highlights complementary strengths of the ShapG and PFI methods.
+[^2] Percentage changes in Table 3 are computed per-dataset and then averaged across all datasets.
+ShapG performs exceptionally well on classification tasks, achieving only an average accuracy reduction of 3.5% while delivering the largest reductions in training time across all models demonstrating its strong efficiency advantage.
+In contrast, the PFI method shows remarkable robustness in regression tasks, with an average drop of only 1.4% in $R^{2}$, making it highly reliable when predictive stability is essential.
+Taken together, these results show that the ShapG method offers superior computational benefits, whereas the PFI method excels in preserving regression performance, providing two complementary and effective strategies for feature selection.
+4 Conclusions and future work
 In this work, we proposed a feature selection pipeline based on game-theoretic feature importance evaluation.
-Our pipeline employs a k-means based diversity sampling procedure to select a representative subset of observations for feature importance evaluation, thereby reducing the influence of outliers and ensuring more robust feature ranking.
-We demonstrated the effectiveness of this approach on 10 benchmark datasets using state-of-the-art tabular deep learning models.
-The results show that integrating our feature selection pipeline into model training reduces computational costs by 12.3% on average while maintaining competitive performance, with accuracy degradation by 5.9% on average across all tasks.
+The pipeline employs a $k$-means-based diversity sampling procedure to obtain a representative subset of observations, thereby reducing the influence of outliers and enabling more stable feature ranking.
+We evaluated two variants of this pipeline, namely, ShapG and PFI approaches, across ten benchmark datasets using state-of-the-art tabular deep-learning models.
+The results show that the proposed pipeline consistently reduces training cost while maintaining competitive predictive performance.
+As summarized in Table 3, ShapG offers the largest reduction in training time with only a modest loss in classification accuracy, whereas the permutation-based method exhibits higher robustness on regression tasks and causes only minimal degradation in the corresponding metrics.
+These observations indicate that ShapG is a suitable choice when computational efficiency is prioritized, while the permutation-based method is preferable when the stability of regression performance is of primary importance.
 Our study has several limitations that open opportunities for future research.
-First, we evaluated only a single feature selection configuration (retaining the top 80% of features by importance) using the ShapG method, which showed effectiveness in recent feature importance evaluation studies Zhao et al.
-(2025a; b).
-Comprehensive comparisons with alternative feature selection approaches—including LIME, Permutation Feature Importance, and other game-theoretic methods with different characteristic functions—remain to be explored.
+First, we evaluated only two feature selection configurations (retaining the top 80% of features by importance) using the ShapG and PFI methods.
+Comprehensive comparisons with alternative feature selection approaches, including LIME and other game-theoretic methods with different characteristic functions, remain to be explored.
 We can also extend the proposed method by varying the number of features selected (the percentage of $q$-top features) for a fixed drop in the model performance.
 This extension is due to varying the number of significant features depending of the dataset and deep-learning model using as a basis.
-Additionally, extending this benchmark to industrial datasets, such as those described in Rubachev et al. (2024), would provide valuable insights on performance of real-world data analysis.
+Additionally, extending this benchmark to industrial datasets, such as those described in rubachev2024tabred, would provide valuable insights on performance of real-world data analysis.
 The proposed diversity sampling method also shows promising results beyond feature importance evaluation.
-Preliminary experiments suggest its effectiveness for training data sampling, in particular, for domains with heterogeneous data distributions.
-While we have observed encouraging results on proprietary wireless communication datasets with TabM (Gorishniy et al., 2025) and FT-transformer (Gorishniy et al., 2021) models, we cannot present these findings due to confidentiality constraints.
-Future work should investigate this application systematically on publicly available datasets.
-These findings suggest that feature selection remains a valuable preprocessing step even for modern deep learning approaches to tabular data, offering a practical way to data analysts to balance computational efficiency with respect to predictive performance.
+Preliminary experiments suggest its effectiveness for training data sampling, in particular, for domains with heterogeneous data distributions, e.g.
+wireless network data reported by base stations (measurement reports).
+While we have observed encouraging results on proprietary wireless communication datasets with TabM (gorishniy2024tabm) and FT-transformer (gorishniy2021revisiting) models, we cannot present these findings due to confidentiality constraints.
+Future work should conduct a systematic comparison between the proposed diversity sampling method and existing cluster-based approach (axiotis2024data), and explore these applications on publicly available datasets.
+These findings suggest that feature selection remains a valuable preprocessing step even for modern deep-learning approaches to tabular data, offering a practical way to data analysts to balance computational efficiency with respect to predictive performance.